If one-half a number is subtracted from five-sixths of the number, the difference is 6. Find the number. (5/6x-1/2x=6)
Answer by jsteppergurl(2)

Well you're going to set up this equation using x as your variable. So you get 5/6x-1/2x=6. However, you also want to get the fractions to have a common denominator. 1/2=?/6 1/2=3/6 so now it is 5/6x-3/6x=6. Now subtract 5-3, so you get 2/6x=6. Simplify 2/6x to 1/3x=6. Finally multiply both sides of the equation by 3 (since 3 is the denominator of the fraction) and you get 1x=18. (1/3*3=1 and 6*3=18.) And there's your answer! x=18